Circular No. 7293

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

HDE 245770 F. Giovannelli, Istituto di Astrofisica Spaziale, CNR, Rome; L. Sabau-

Graziati, División de Ciencias del Espacio, INTA, Madrid; S. Bernabei, Osservatorio Astronomico di Bologna; and S. Galleti, Dipartimento di Astronomia, Universitá di Bologna, write: “Low-resolution spectrosopy (Δλ ∼ 0.66 nm) of HDE 245770 (the optical counterpart of the x-ray pulsar A0535+26; e.g., IAUC 3167, 3655, 5746; Giovannelli and Sabau-Graziati 1992, Space Sci. Rev. 59, 1) made on Oct. 8 with the 1.5-m Loiano telescope reveals Hα in emission with EW ∼ −0.5 nm. Also, He I lines (587.6, 667.8, and 706.5 nm) are seen in emission and show doublings separated by ∼ 1.5 nm. Hβ is in weak absorption (EW ∼ +0.017 nm). Hγ and Hδ are seen in absorption (EW ∼ +0.16 and +0.30 nm, respectively). This result clearly shows the resumption of optical activity after a period from early 1996 to early 1999 in which a depletion of the circumstellar disk was clearly seen (Piccioni et al. 1999, IAU Coll. No. 175, Alicante, Spain). For instance, Hα was in absorption on 1998 Nov. 11 and 12 with EW ∼ +0.04 and +0.05 nm, respectively. Also Hβ, Hγ, and Hδ were in absorption on 1998 Nov. 9 with EW = +0.19, +0.28, and +0.34 nm, respectively. Photometry of HDE 245770 performed on 1999 Oct. 9 and 11 gives B = 9.78 ± 0.04 and 9.77 ± 0.04, respectively. The historical B lightcurve has in the past 15 yr shown a continuous trend toward lower luminosity (starting from a stable mag of ∼ 9.5 during the mid-1970s to the mid-1980s, and fading to mag ∼ 10), but this trend is now reversing. As a consequence of this behavior, we expect the x-ray pulsar companion A0535+26 to have a normal x-ray outburst (following the classification of Giovannelli and Sabau-Graziati, ibid.) of intensity ∼ 0.1 Crab on 1999 Dec. 27–31. Optical indicators (e.g., Hγ, usually in absorption, will enhance to fill the continuum or even show emission; optical UBVRI flare of intensity ∼ 0.1 mag) will precede the x-ray outburst by ∼ 7–10 days. This prediction is

based on RXTE data (http://space.mit.edu/XTE/asmlc/ASM.htlm) and

on the orbital period of the system (110.8 days, from the best-fit multi-year optical data of Giovannelli’s group and others; see Guarnieri et al. and de Martino et al. 1985, in Multifrequency Behaviour of Galactic Accreting Sources, pp. 310 and 326; and Priedhorsky and Terrell 1983, Nature 303, 681, who derived a period of 111.0 ± 0.4 days from VELA-5B data, related to x-ray quiescent emission from A0535+26).”

1999 October 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green